4b Covert into rectangular coordinates (4,-2π÷3)

1
Expert's answer
2022-06-09T11:02:09-0400
"x=4\cos(-\dfrac{2\pi}{3})=-2"

"y=4\sin(-\dfrac{2\pi}{3})=-2\sqrt{3}"

"(-2, -2\sqrt{3})"